### v6

@@ 1 Introduction
In spite of the growing dominance of deep neural networks, modeling and learning of linear dynamical systems (LDS) remain relevant due to the availability of analytical solutions, significantly lower computational burden, and ease of analyses and inspection. LDS continue to find use in diverse applications such as computer vision (Chan and Vasconcelos, 2005; Boots et al., 2007; Chien et al., 2022; Wang and Mu, 2024) and time series prediction (Liu et al., 2024; Takeishi et al., 2017). Even if the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory (Koopman, 1931; Williams et al., 2015; Mezic, 2020) to approximate nonlinear systems with linear systems in a higher-dimensional lifted state space. These techniques further increase the relevance and importance of learning LDS in complex applications, such as legged robotics (Kim et al., 2024; Li et al., 2024), robotic manipulation (Bruder et al., 2021; Han et al., 2023, 2024; Chen et al., 2024), atmospheric sciences (Navarra et al., 2021; Millard et al., 2024; Zheng et al., 2022), and fluid dynamics (Mezić, 2013; Schmid, 2022).
When time-series data associated with system evolution is available, learning LDS can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective (Ljung, ). However, such unconstrained optimization methods tend to ignore important properties of the underlying systems, such as stability. Consequently, even if the underlying system is known to be stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
A number of methods have been developed to enforce stability of learned LDS and reduce sensitivity to training data (Lacy and Bernstein, 2002, 2003; Boots et al., 2007; Huang et al., 2016; Gillis et al., 2020; Mamakoukas et al., 2020, 2023). These methods take one of three approaches: a) minimizing the regression objective under stability constraints (Lacy and Bernstein, 2002, 2003), b) iterating between solving an unconstrained problem and refining the solution under stability constraints (Boots et al., 2007; Huang et al., 2016), and c) iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (Gillis et al., 2020; Mamakoukas et al., 2020, 2023). In general, these methods inevitably introduce a considerable computational burden, potentially countering one of the primary advantages of linear techniques.
@@
In contrast to existing approaches discussed above, SC offers stability-by-construction without requiring constrained optimization or iterative updates. SC can efficiently scale to high-dimensional systems (e.g., image prediction). Further, seminal works on the Koopman operator (Rowley et al., 2009; Mezić, 2013; Mezic, 2020) reveal fundamental connections between the spectrum of the Koopman matrix and desirable properties in learned nonlinear systems. We show that SC can be combined with Koopman operators to enhance the predictive accuracy and stability of learned nonlinear dynamical systems (e.g., robotic systems).
While it can be readily seen that SC ensures stability by directly manipulating the eigenvalues, it might seem puzzling that it does not hurt the model’s predictive accuracy. To explain SC’s surprising effectiveness, we provide practical intuitions and theoretical insights (e.g., see Fig. 1 for a geometric interpretation of SC). Indeed, prior works have identified several linear techniques that defy expectations: controlling the singular values of linear layers improves training stability and generalization (Boroojeny et al., 2024; Sedghi et al., 2018; Senderovich et al., 2022); retraining only the last linear layer improves robustness to spurious correlations (Kirichenko et al., 2022); and Low-Rank Adaptation (LoRA) dramatically improves efficiency (Hu et al., 2022).
We conducted comprehensive experiments involving benchmark datasets across different domains (two video prediction datasets (Chan and Vasconcelos, 2005; Hadji and Wildes, 2018), and a robot model learning dataset (Gaz et al., 2019)) and strong baselines (CG (Boots et al., 2007), WLS (Huang et al., 2016), and SOC (Mamakoukas et al., 2020)). Our results reveal that SC consistently achieves comparable (if not better) predictive accuracy, and generates stable long-horizon predictions. Notably, SC is able to achieve this level of performance while being orders-of-magnitude faster.
Our experiments also show that SC can be readily combined with Koopman operators to learn nonlinear systems capable of stable and consistent predictions in mere seconds (e.g., behavioral dynamics of dexterous manipulation skills for a multi-fingered robotic hand (Han et al., 2023)). Importantly, we show that SC enables the learned system to generate robust and stable predictions even when the training data includes unsuccessful or truncated demonstrations (commonly characteristics of real-world robot learning datasets (Qin et al., 2022; Shaw et al., 2023; Bahl et al., 2022)). In contrast, we find unconstrained learning yields unstable LDS leading to erratic and ineffective robot behaviors in the presence of such challenging characteristics.
@@ 2 Background: Learning Stable Linear Dynamical Systems
Unconstrained Learning: In its simplest form, the problem of learning an LDS from $D$ reduces to solving a standard least-squares problem. (Montgomery et al., 2021; Seber and Lee, 2012):
@@
While solving the optimization problem in Eq. (2) can yield the constrained optimal solution which is guaranteed to be stable, it is known to be non-convex and requires approximations (Boyd and Vandenberghe, 2004; Mamakoukas et al., 2020).
@@ 3 Post-Hoc Stability via Spectral Clipping
Stability: The reconstructed linear system after the above clipping procedure, $x_{t+1}=\hat{A}_{SC}\ x_{t}$ can be trivially shown to be asymptotically (marginally) stable When $\varepsilon>0$ ($\varepsilon=0$) since the maximum eigenvalue of $\hat{A}$ is strictly less than (equal to) one by construction (Robinson, 2012).
@@ 4 A Koopman-based Adaptation of SC to Nonlinear Systems
We also investigate the potential utility of SC when learning nonlinear dynamical systems (NLDS) by combining it Koopman operator-based techniques (Koopman, 1931; Mezic, 2020).
Let $\xi_{k+1}=f(\xi_{k})$ be an unknown discrete-time NLDS with $\xi_{k}\in\Xi\subseteq\mathbb{R}^{n},\forall k\in\mathbb{N}$ and $D_{f}=\{[\xi_{1}^{(j)},\xi_{2}^{(j)},\cdots,\xi_{T}^{(j)}]\}_{j=1}^{N_{f}}$ be a dataset containing $N_{f}$ trajectories of the NLDS. Koopman operator-based learning approaches learn to approximately encode the NLDS as $\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$, where $\hat{K}\in\mathbb{R}^{m\times m}$ is the learned finite-dimensional approximation of the Koopman operator $\mathrm{z}_{k}=\phi(\xi_{k})\in\mathbb{R}^{m}$ is the “lifted" latent state produced by the so-called “lifting" function $\phi:\Xi\rightarrow\mathcal{Z}$ with $m\gg n$, and $\mathcal{Z}$ is the approximated Koopman-invariant subspace in which the dynamics appear linear (Koopman, 1931; Mezic, 2020).
While it is possible to learn or design an encoder and a decoder to lift and reconstruct the original state, other formulations allow for trivial decoding and enable further analysis. In particular, we investigate state-inclusive latent space designs (Proctor et al., 2018; Han et al., 2023; Johnson et al., 2025) which append the original state to the nonlinear lifted embedding (i.e., $\mathrm{z_{t}}=[\phi(\xi)^{T},\xi^{T}]^{T}$ with $\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$). As such, a trivial linear map can recover the original states from the latent states.
@@
Open-loop predictive stability: Indeed, stability in the latent space does not necessarily imply stability in the original state space of the NLDS, and in fact requires assumptions about forward and backward conjugacy Lutkus et al. (2025). As such, we do not claim that SC ensures closed-loop stability of the underlying NLDS. However, SC ensures that lifted latent LDS is stable by construction in $\mathcal{Z}$. Given any initial condition ($\mathrm{z_{0}}=[\phi(\xi_{0})^{T},\xi_{0}^{T}]^{T}$), the predictions generated by the latent linear system is guaranteed to converge to the origin ($\hat{\mathrm{z}}_{k}\rightarrow 0$ as $k\rightarrow\infty$). Since we employ state-inclusive lifting ($\mathrm{\hat{z}_{k}}=[\phi(\hat{\xi}_{k})^{T},\hat{\xi}_{k}^{T}]^{T}$), the predicted original states $\hat{\xi}$ are merely a subset of the predicted latent states $\hat{\mathrm{z}}_{k}$ and as such will also converge to the origin ($\hat{\xi}_{k}\rightarrow 0$ as $k\rightarrow\infty$). We refer to this property as open-loop predictive stability.
@@ 5.1 Experimental Design
Baselines: We compared SC against the following baselines: i) LS: vanilla least squares, ii) (CG): constraint generation (Boots et al., 2007), iii) WLS: weighted least squares (Huang et al., 2016), and iv) SOC: a characterization approach (Mamakoukas et al., 2020, 2023).
@@
i) UCSD (Chan and Vasconcelos, 2005): This dataset contains 254 highway traffic videos, each consisting of 48–52 frames of size $48\times 48$. Following Mamakoukas et al. (2020). we apply SVD to each sequence and retain subspaces of dimension $r\in\{3,30\}$ and reduce dimensionality.
ii) DTDB (Hadji and Wildes, 2018): A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying number of frames. As we did for the UCSD dataset, we leveraged SVD to lower the state dimension to $r=300$.
iii) DexManip (Han et al., 2023; Rajeswaran et al., 2018b): A robotic dexterous manipulation dataset containing demonstrations of a 23-DoF multi-fingered Adroit hand performing four complex tasks in MUJOCO (Todorov et al., 2012).
iv) Franka Panda (Gaz et al., 2019): A dataset of Franka Panda Robot Arm’s movements as a consequence of different control inputs. Results are reported in Appendix E.
@@ 5.4 SC enables consistent long-horizon predictions
To quantify long-term predictive performance, we use the following metrics: i) NIQE quality score (Mittal et al., 2012)): A common image quality metric assessing the meaningfulness of LDS predictions. Note that some poor predictions (e.g., when most pixel values blow up or converge to 0) can produce unbounded NIQE scores, so we substitute them with 10000; ii) Moving Ratio  [^1] Moving is defined as $F_{M}-F_{M-2}||_{1}>T$, where $\{F_{i}\}_{i=1}^{M}$ is a sequence of $M$ image frames and $T=9$.: The ratio of sequences whose final frames still exhibit the dynamic pattern, indicating whether LDS evolution preserves it.
@@ 5.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
We evaluated the effectiveness of combining the Koopman Operator theory (Koopman, 1931), SC can be effectively and efficiently applied to highly nonlinear robot manipulation tasks by evaluating SC on DexManip (Rajeswaran et al., 2018b), a robotic dexterous manipulation dataset containing demonstrations of four complex tasks using a 23-DoF multi-fingered hand. Though the original state evolution is nonlinear, it was recently demonstrated that the evolution can be encoded using a linear system when “lifted" to a higher-dimensional space (e.g., $r\approx 750$) that approximates the Koopman invariant subspace Han et al. (2023). Four tasks are briefly described as follows: i) Pen Reorientation: The robot hand is tasked with orienting a pen to a random target rotation angles, ii) Door Opening: The robot hand is tasked with opening a door from a random starting position, iii) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and iv) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. F for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
@@ 5.6 SC offers robustness to unsuccessful and shorter demonstrations
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in Rowley et al. (2009), we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset. See detailed experiments in Appendix. D, where we demonstrated how SC effectively stabilizes unstable modes and keeps stable modes.
@@ Appendix A Approximation error bounds for non-diagonalizable system matrices
Lemma. A.1 can be proved by the fact that the set of real matrices that are diagonalizable over $\mathbb{C}$ is dense in $\mathbb{R}^{n\times n}$. The details can be found from the matrix perturbation theory (Horn and Johnson, 2012; Trefethen and Bau, 2022).
Theorem A.3.
For any $\eta>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, we can always find its diagonalizable counterpart $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.
@@
$\tilde{A}_{LS}$ can be denoted as
$\tilde{A}_{LS}=\hat{A}_{LS}+E$ (3)
where $E\in\mathbb{R}^{n\times n}$ is the residual matrix and, from Lemma. A.1, $E||_{1}$ can be arbitrarily small (i.e., $E||_{1}<\gamma$). Then we have
$\tilde{A}_{LS}^{k}\mathbf{z}=(\hat{A}_{LS}+E)^{k}\mathbf{z}=\hat{A}_{LS}^{k}\mathbf{z}+\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}$ (4)
Therefore,
$\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}=-\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}$ (5)
Using triangle inequality, we have
[…]
Since $\hat{A}_{LS}$, $k$ and $\mathbf{z}$ are all bounded, there exists a bounded $m>0$, s.t.
$\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<m\gamma$ (7)
Let $\eta=m\gamma$. Note that since Lemma. A.1 guarantees the existence of a diagonalizable matrix $\tilde{A}_{LS}$ that satisfies an arbitrarily small $\gamma$, it follows that $\eta=m\gamma$ can also be made arbitrarily small. Therefore, for any $\eta>0$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.

### v7

@@ 1 Introduction
In spite of the growing dominance of deep neural networks, modeling and learning of linear dynamical systems (LDS) remain relevant due to the availability of analytical solutions, significantly lower computational burden, and ease of analyses and inspection. LDS continue to find use in diverse applications such as computer vision (chan2005probabilistic; boots2007constraint; chien2022gpu; wang2024localized) and time series prediction (liu2024koopa; takeishi2017learning). Even if the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory (Koopman1931Koopman; williams2015data; mezic2020koopman) to approximate nonlinear systems with linear systems in a higher-dimensional lifted state space. These techniques further increase the relevance and importance of learning LDS in complex applications, such as legged robotics (kim2024learning; li2024continual), robotic manipulation (Bruder2021Soft; han2023KODex; han2024learning; chen2024korol), atmospheric sciences (navarra2021estimation; millard2024deep; zheng2022hybrid), and fluid dynamics (mezic2013analysis; schmid2022dynamic).
When time-series data associated with system evolution is available, learning LDS can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective (ljung1998system). However, such unconstrained optimization methods tend to ignore important properties of the underlying systems, such as stability. Consequently, even if the underlying system is known to be stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
A number of methods have been developed to enforce stability of learned LDS and reduce sensitivity to training data (lacy2003subspace_conf; lacy2003subspace; boots2007constraint; huang2016learning; gillis2020note; mamakoukas2020memory; mamakoukas2023learning). These methods take one of three approaches: a) minimizing the regression objective under stability constraints (lacy2003subspace_conf; lacy2003subspace), b) iterating between solving an unconstrained problem and refining the solution under stability constraints (boots2007constraint; huang2016learning), and c) iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (gillis2020note; mamakoukas2020memory; mamakoukas2023learning). In general, these methods inevitably introduce a considerable computational burden, potentially countering one of the primary advantages of linear techniques.
@@
In contrast to existing approaches discussed above, SC offers stability-by-construction without requiring constrained optimization or iterative updates. SC can efficiently scale to high-dimensional systems (e.g., image prediction). Further, seminal works on the Koopman operator (rowley2009spectral; mezic2013analysis; mezic2020koopman) reveal fundamental connections between the spectrum of the Koopman matrix and desirable properties in learned nonlinear systems. We show that SC can be combined with Koopman operators to enhance the predictive accuracy and stability of learned nonlinear dynamical systems (e.g., robotic systems).
While it can be readily seen that SC ensures stability by directly manipulating the eigenvalues, it might seem puzzling that it does not hurt the model’s predictive accuracy. To explain SC’s surprising effectiveness, we provide practical intuitions and theoretical insights (e.g., see Fig. 1 for a geometric interpretation of SC). Indeed, prior works have identified several linear techniques that defy expectations: controlling the singular values of linear layers improves training stability and generalization (boroojeny2024spectrum; sedghi2018singular; senderovich2022towards); retraining only the last linear layer improves robustness to spurious correlations (kirichenko2022last); and Low-Rank Adaptation (LoRA) dramatically improves efficiency (hu2022lora).
We conducted comprehensive experiments involving benchmark datasets across different domains (two video prediction datasets (chan2005probabilistic; hadji2018new), and a robot model learning dataset (gaz2019dynamic)) and strong baselines (CG (boots2007constraint), WLS (huang2016learning), and SOC (mamakoukas2020memory)). Our results reveal that SC consistently achieves comparable (if not better) predictive accuracy, and generates stable long-horizon predictions. Notably, SC is able to achieve this level of performance while being orders-of-magnitude faster.
Our experiments also show that SC can be readily combined with Koopman operators to learn nonlinear systems capable of stable and consistent predictions in mere seconds (e.g., behavioral dynamics of dexterous manipulation skills for a multi-fingered robotic hand (han2023KODex)). Importantly, we show that SC enables the learned system to generate robust and stable predictions even when the training data includes unsuccessful or truncated demonstrations (commonly characteristics of real-world robot learning datasets (qin2022dexmv; shaw2023videodex; bahl2022human)). In contrast, we find unconstrained learning yields unstable LDS leading to erratic and ineffective robot behaviors in the presence of such challenging characteristics.
@@ 2 Background: Learning Stable Linear Dynamical Systems
Unconstrained Learning: In its simplest form, the problem of learning an LDS from $D$ reduces to solving a standard least-squares problem. (montgomery2021introduction; seber2012linear):
@@
While solving the optimization problem in Eq. (2) can yield the constrained optimal solution which is guaranteed to be stable, it is known to be non-convex and requires approximations (boyd2004convex; mamakoukas2020memory).
@@ 3 Post-Hoc Stability via Spectral Clipping
Stability: The reconstructed linear system after the above clipping procedure, $x_{t+1}=\hat{A}_{SC}\ x_{t}$ can be trivially shown to be asymptotically (marginally) stable When $\varepsilon>0$ ($\varepsilon=0$) since the maximum eigenvalue of $\hat{A}$ is strictly less than (equal to) one by construction (robinson2012introduction).
@@ 4 A Koopman-based Adaptation of SC to Nonlinear Systems
We also investigate the potential utility of SC when learning nonlinear dynamical systems (NLDS) by combining it with Koopman operator-based techniques (Koopman1931Koopman; mezic2020koopman).
Let $\xi_{k+1}=f(\xi_{k})$ be an unknown discrete-time NLDS with $\xi_{k}\in\Xi\subseteq\mathbb{R}^{n},\forall k\in\mathbb{N}$ and $D_{f}=\{[\xi_{1}^{(j)},\xi_{2}^{(j)},\cdots,\xi_{T}^{(j)}]\}_{j=1}^{N_{f}}$ be a dataset containing $N_{f}$ trajectories of the NLDS. Koopman operator-based learning approaches learn to approximately encode the NLDS as $\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$, where $\hat{K}\in\mathbb{R}^{m\times m}$ is the learned finite-dimensional approximation of the Koopman operator. $\mathrm{z}_{k}=\phi(\xi_{k})\in\mathbb{R}^{m}$ is the “lifted" latent state produced by the so-called “lifting" function $\phi:\Xi\rightarrow\mathcal{Z}$ with $m\gg n$, and $\mathcal{Z}$ is the approximated Koopman-invariant subspace in which the dynamics appear linear (Koopman1931Koopman; mezic2020koopman).
While it is possible to learn or design an encoder and a decoder to lift and reconstruct the original state, other formulations allow for trivial decoding and enable further analysis. In particular, we investigate state-inclusive latent space designs (proctor2018generalizing; han2023KODex; johnson2025heterogeneous) which append the original state to the nonlinear lifted embedding (i.e., $\mathrm{z_{t}}=[\phi(\xi_{t})^{T},\xi_{t}^{T}]^{T}$ with $\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$). As such, a trivial linear map can recover the original states from the latent states.
@@
Open-loop predictive stability Clipping ensures that the learned latent linear predictor $z_{t+1}=K_{\mathrm{SC}}z_{t}$ satisfies $\rho(K_{\mathrm{SC}})\leq 1-\epsilon$ for some $\epsilon\geq 0$. If $\epsilon>0$, then $K_{\mathrm{SC}}^{t}z_{0}\to 0$ as $t\to\infty$ for all $z_{0}$, so the latent predictions remain bounded and decay to the origin. If $\epsilon=0$, then $\rho(K_{\mathrm{SC}})\leq 1$ implies boundedness only under additional conditions (e.g., no defective Jordan blocks on the unit circle), and in general one can only guarantee that the latent predictor is not exponentially unstable. Since we employ a state-inclusive lifting (i.e., the original state appears as a subvector of $z$), the decoded open-loop predictions inherit the same boundedness/decay properties as the latent trajectory. We emphasize that this is a statement about the learned open-loop predictor, and does not imply closed-loop stability of the underlying nonlinear dynamical system without further assumptions.
@@ 5.1 Experimental Design
Baselines: We compared SC against the following baselines: i) LS: vanilla least squares, ii) (CG): constraint generation (boots2007constraint), iii) WLS: weighted least squares (huang2016learning), and iv) SOC: a characterization approach (mamakoukas2020memory; mamakoukas2023learning).
@@
i) UCSD (chan2005probabilistic): This dataset contains 254 highway traffic videos, each consisting of 48–52 frames of size $48\times 48$. Following mamakoukas2020memory. we apply SVD to each sequence and retain subspaces of dimension $r\in\{3,30\}$ and reduce dimensionality.
ii) DTDB (hadji2018new): A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying number of frames. As we did for the UCSD dataset, we leveraged SVD to lower the state dimension to $r=300$.
iii) DexManip (han2023KODex; Rajeswaran2018DAPG): A robotic dexterous manipulation dataset containing demonstrations of a 23-DoF multi-fingered Adroit hand performing four complex tasks in MUJOCO (Todorov2012MUJuCO).
iv) Franka Panda (gaz2019dynamic): A dataset of Franka Panda Robot Arm’s movements as a consequence of different control inputs. Results are reported in Appendix E.
@@ 5.4 SC enables consistent long-horizon predictions
To quantify long-term predictive performance, we use the following metrics: i) NIQE quality score (mittal2012making)): A common image quality metric assessing the meaningfulness of LDS predictions. Note that some poor predictions (e.g., when most pixel values blow up or converge to 0) can produce unbounded NIQE scores, so we substitute them with 10000; ii) Moving Ratio  [^1] Moving is defined as $F_{M}-F_{M-2}||_{1}>T$, where $\{F_{i}\}_{i=1}^{M}$ is a sequence of $M$ image frames and $T=9$.: The ratio of sequences whose final frames still exhibit the dynamic pattern, indicating whether LDS evolution preserves it.
@@ 5.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
We evaluated the effectiveness of combining the Koopman Operator theory (Koopman1931Koopman), SC can be effectively and efficiently applied to highly nonlinear robot manipulation tasks by evaluating SC on DexManip (Rajeswaran2018DAPG), a robotic dexterous manipulation dataset containing demonstrations of four complex tasks using a 23-DoF multi-fingered hand. Though the original state evolution is nonlinear, it was recently demonstrated that the evolution can be encoded using a linear system when “lifted" to a higher-dimensional space (e.g., $r\approx 750$) that approximates the Koopman invariant subspace han2023KODex. Four tasks are briefly described as follows: i) Pen Reorientation: The robot hand is tasked with orienting a pen to a random target rotation angles, ii) Door Opening: The robot hand is tasked with opening a door from a random starting position, iii) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and iv) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. F for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
@@ 5.6 SC offers robustness to unsuccessful and shorter demonstrations
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in rowley2009spectral, we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset. See detailed experiments in Appendix. D, where we demonstrated how SC effectively stabilizes unstable modes and keeps stable modes.
@@ Appendix A Approximation error bounds for non-diagonalizable system matrices
Lemma. A.1 can be proved by the fact that the set of real matrices that are diagonalizable over $\mathbb{C}$ is dense in $\mathbb{R}^{n\times n}$. The details can be found from the matrix perturbation theory (horn2012matrix; trefethen2022numerical).
Theorem A.3 (Finite-horizon approximation of trajectories).
Let $K\in\mathbb{N}$ represent a fixed finite horizon of interest. For any $\varepsilon>0$, any $A\in\mathbb{R}^{n\times n}$ (possibly non-diagonalizable), and any $z\in\mathbb{R}^{n}$, there exists a matrix $\tilde{A}\in\mathbb{R}^{n\times n}$ that is diagonalizable over $\mathbb{C}$ such that for all $k\in\{0,1,\dots,K\}$,
$\|A^{k}z-\tilde{A}^{k}z\|_{2}\leq\varepsilon.$
@@
Let $E:=\tilde{A}-A$. For any integer $k\geq 1$, the following identity holds without any commutativity assumptions:
$\tilde{A}^{k}-A^{k}=\sum_{i=0}^{k-1}\tilde{A}^{\,k-1-i}\,E\,A^{i}.$
Therefore, by submultiplicativity of the operator norm,
$\|\tilde{A}^{k}-A^{k}\|_{2}\leq\|E\|_{2}\sum_{i=0}^{k-1}\|\tilde{A}\|_{2}^{k-1-i}\|A\|_{2}^{i}.$
Choose $\|E\|_{2}\leq 1$, which implies $\|\tilde{A}\|_{2}\leq\|A\|_{2}+\|E\|_{2}\leq\|A\|_{2}+1=:\alpha$. Then for all $k\leq K$,
$\|\tilde{A}^{k}-A^{k}\|_{2}\leq\|E\|_{2}\sum_{i=0}^{k-1}\alpha^{k-1-i}\alpha^{i}=\|E\|_{2}\,k\,\alpha^{k-1}\leq\|E\|_{2}\,K\,\alpha^{K-1}.$
Hence,
$\|A^{k}z-\tilde{A}^{k}z\|_{2}\leq\|\tilde{A}^{k}-A^{k}\|_{2}\|z\|_{2}\leq\|E\|_{2}\,K\,\alpha^{K-1}\,\|z\|_{2}.$
Now choose
$\|E\|_{2}\leq\min\Bigl\{1,\ \frac{\varepsilon}{K\,\alpha^{K-1}\,\|z\|_{2}}\Bigr\}.$
Given the density of diagonalizable matrices in $\mathbb{R}^{n\times n}$ (see Lemma 1), we can achieve the bound in $E$ with an appropriate diagonalizable $\tilde{A}$. This yields the desired bound for all $k\in\{0,\dots,K\}$.
Remark. The above theorem is specifically stated for a fixed finite horizon $K$. Without additional assumptions (e.g., power-boundedness of $\hat{A}_{LS}$ and $\tilde{A}_{LS}$), a uniform bound as $k\to\infty$ generally cannot be guaranteed since $\|A^{k}\|$ may grow with $k$.